OpenStudy (anonymous):

hi guys , i am troubled with this question the 8th term of an arithmetic sequence is 25 and the 20th term is -11. find the first term of the sequence and the common difference and use these to write the general term

OpenStudy (amistre64):

you know the 8th and the 20 th terms; 20 - 8 = 12 terms between them, 25 + 12*d = -11 if i see it right;

OpenStudy (amistre64):

solve for d to determine the "common difference"

OpenStudy (amistre64):

then you would simply resolve for a{1} a{8} = a{1} + 8(d) ; they give you a{8} = 25; so solve it for a{1} :)

OpenStudy (anonymous):

how to solve for common difference? i really dont get this problem buddy

OpenStudy (anonymous):

\[t_8=a+7d=25\] \[t_{20}=a+19d=-11\] \[t_{20}-t_8=12d=-36\] \[d=-3\] a+7d=25 a-21=25 a=46 the general term \[t_n=46+((n-1)-3)\] \[t_n=46-3n+3\] \[t_n=49-3n\]

OpenStudy (amistre64):

there is a common difference that keeps adding/subtracting to the problem with each iteration; spose we start at a{8} and treat it as the first term; then a{n} = a{n-1} + d ; 12 times to get to -11 a{12} = a{1} + 12d -11 = 25 + 12d ; subtract out the 25 -36 = 12d ; now divide out the 12 -3 = d

OpenStudy (amistre64):

the common difference; if I did it right is -3

OpenStudy (amistre64):

Now they want you to solve for the first term; but we already know a{8} = 25 a{8} = a{1} +8(-3) 25 = a{1} - 24 ; add 24 25 + 24 = a{1} 49 = a{1} .. maybe

OpenStudy (anonymous):

thank u guys. u are a real life saver